OpenStudy (anonymous):

A dance instructor charges each student $57 for dance lessons. Students who register early receive a $15 discount. A total of 11 students took dance lessons. If 5 of these students registered early,how much should the instructor collect from these 11 students? A $462 B $552 C $342 D $537

OpenStudy (ashleyisakitty):

57-15=42 42*5=210 6*57=342 342+210=552

OpenStudy (texaschic101):

I agree :)

OpenStudy (anonymous):

thank